# Instruction-tuning Aligns LLMs to the Human Brain

### v1

@@ 4.2 Factors underlying LLM-brain alignment
To identify factors underlying the representational similarity between LLMs and human brains, we compute the Pearson correlation between LLM brain alignment and various properties of LLMs: performance on a benchmark involving different reasoning abilities (BBH benchmark; Suzgun et al., 2022), performance on a benchmark requiring domain-specific world knowledge (MMLU; Hendrycks et al., 2021), language modeling ability, and model size.
@@
World Knowledge and Reasoning abilities   MMLU is designed to measure the world knowledge of LLMs across many subject domains. It contains 57 tasks, categorized by the subject domain of world knowledge tested: STEM, Humanities, Social Sciences, and Others (a broad category ranging from finance to marketing to professional medicine). BBH contains 23 tasks, grouped into four categories based on the type of problem-solving ability tested: Algorithmic and Multi-Step Arithmetic Reasoning; Natural Language Understanding; Use of World Knowledge; and Multilingual Knowledge and Reasoning. For both benchmarks, we use category classifications from the original papers. We measure the performance of LLMs on BBH and MMLU using the instruct-eval repository [^2] https://github.com/declare-lab/instruct-eval with default settings (3-shots, 5-shots respectively) and preset prompts. We measure the Pearson correlation (and its p-value) between LLM brain alignment and performance on each category of the MMLU and BBH benchmarks. We obtain p-values after performing false discovery rate correction, ensuring our results are significant and not due to random chance.
@@
We find that brain alignment is significantly and strongly correlated with world knowledge. On the MMLU benchmark, we observe a high correlation between brain alignment scores and the MMLU Overall score (r = 0.81), which reports the mean performance across all world knowledge subject domains on MMLU (Figure 2A). Similarly, brain alignment is also significantly and strongly correlated with the mean performance on tasks included in the world knowledge category of the BBH benchmark (r = 0.68; Figure 2B). Interestingly, we do not find strong correlations with other dimensions of the BBH benchmark (e.g., Algorithmic reasoning and Multilingual reasoning, see Table 2), though this could also be due to limitations of the tested models (most are primarily pretrained on English language data), as indicated by their low raw performance scores on some tasks. Overall, our results provide a strong signal that more accessible representations of world knowledge are a key factor in aligning LLM representations to human brain activity.
@@
Language Modeling Ability   Prior works have shown correlations between brain alignment and next-word prediction (NWP) ability (Caucheteux & King, 2022; Schrimpf et al., 2021). We find similar results for correlation between brain alignment and NWP loss (r = -0.54, Appendix G). Interestingly, the strength of correlation between brain alignment and world knowledge performance (r = 0.81) is greater than that between brain alignment and NWP loss (r = -0.54). This result suggests that world knowledge understanding is a better predictor of brain alignment than NWP ability.
Model Size   Finally, we find that brain alignment is significantly and strongly correlated with model size (r = 0.95), as measured by the logarithm of the number of model parameters (Figure 2C). Schrimpf et al. (2021) observe such a pattern for language models, and we find the pattern holds for instruction-tuned models, and models trained at a larger scale than their study (7B+ parameters).
@@
However, model size alone does not determine brain alignment. Our results show that smaller instruction-tuned LLMs can have greater brain alignment than larger vanilla models. For example, LLaMA-13B obtains brain alignment of 0.220, Vicuna-13B obtains 0.229, LLaMA-33B obtains 0.227, and Vicuna-33B obtains 0.232. Hence, Vicuna-13B has greater brain alignment than LLaMA-33B despite being less than 40% its size. We observe a similar trend in another four models: T5-base, Flan-T5-base, T5-large, Flan-T5-large. Also, prior works have shown that large random models achieve poor brain alignment (Schrimpf et al., 2021). These results demonstrate that there are LLM properties aside from model size that contribute significantly to brain alignment.
@@
Datasets   We observe significant correlations between brain alignment and world knowledge and model size on Pereira2018 and Wehbe2014, but surprisingly not Blank2014. We hypothesize that it may be because Blank2014 has a small participant size (N = 5) that leads to greater noise in the results, compared to Pereira2018 and Wehbe2014.
Table 2: Brain alignment strongly correlates with world knowledge across all subject domains in MMLU, as well as the world knowledge category in BBH. At the same time, brain alignment is not significantly correlated with all other types of problem-solving abilities in BBH (e.g., algorithmic or multilingual reasoning). We obtain p-values after performing false discovery rate correction, ensuring our results are significant and not due to random chance.
Task category Brain Alignment corrected Number Average Model
Correlation (r) p-value of tasks Performance
MMLU – Overall Score 0.809 0.000329 57 0.36
MMLU – STEM 0.792 0.000343 18 0.28
MMLU – Humanities 0.791 0.000343 13 0.34
MMLU – Social Sciences 0.807 0.000329 12 0.41
MMLU – Others 0.809 0.000329 14 0.40
BBH – Overall score 0.384 0.177 23 0.28
BBH – Algorithmic reasoning 0.194 0.558 8 0.22
BBH – Language understanding 0.163 0.585 3 0.43
BBH – World knowledge 0.679 0.005 5 0.36
BBH – Multilingual reasoning -0.035 0.895 1 0.19
BBH – Others 0.478 0.083 6 0.27
@@ 5 Behavioral Alignment
Figure 3: Instruction-tuning LLMs generally does not improve behavioral alignment to human reading times. Furthermore, behavioral alignment correlates poorly with all other tested measures: world knowledge, model size, and next-word prediction (NWP) ability. (A) We present the same language stimuli to LLMs and human participants, using the Futrell2018 benchmark in Brain-Score, which contains naturalistic stories. We compute the behavioral alignment as the Pearson correlation between LLM perplexity for each word and human reading times for each word. (B) Instruction-tuning does not generally improve behavioral alignment. Furthermore, behavioral alignment is poorly and not significantly correlated with all other measures: (C) world knowledge on MMLU (r = 0.08, p = 0.76), (D) model size (r = 0.26, p = 0.31), (E) NWP loss for T5 models (r = -0.2, p = 0.54), and (F) NWP loss for LLaMA models (r = 0.68, p = 0.21).
@@
In the previous section, we show that instruction-tuning aligns the internal representations of LLMs to human brain recordings (Section 4.1). In this section, we explore whether instruction-tuning also aligns LLM behavior to human behavior.
@@
Following the approach previously proposed by Schrimpf et al. (2021) and implemented in the Brain-Score package (Schrimpf et al., 2020), we measure behavioral alignment by evaluating the similarity between LLM per-word perplexity and human per-word reading times, given the same language stimuli (Figure 3A). We use the self-paced reading times dataset from Futrell et al. (2018), consisting of the reading times of 179 human participants recorded while they were visually presented with 10 naturalistic stories. We provide language stimuli from this data as input to LLMs and measure their per-word perplexity. Finally, we evaluate behavioral alignment by computing the Pearson correlation between per-word LLM perplexity and per-word human reading times.
@@
Using the same set of models as in the previous section, we compare the behavioral alignment of each instruction-tuned LLM against its vanilla counterpart. Our results generally indicate that instruction-tuning LLMs does not improve behavioral alignment to human reading times (Figure 3B). For half of the LLMs tested, it results in no change or reduced behavioral alignment. Then, we test the correlation between LLM behavioral alignment and model size, next-word prediction ability, various reasoning abilities (as measured by performance on the BBH benchmark), and world knowledge across various domains (as measured by performance on the MMLU benchmark). Contrary to our findings on the correlation between brain alignment and model size and world knowledge (Section 4.2), we do not find that these factors are correlated with the behavioral alignment of models: world knowledge in Figure 3C (r = 0.08, p = 0.76), model size in Figure 3D (r = 0.26, p = 0.31), next-word prediction loss for T5 models in Figure 3E (r = -0.2, p = 0.54), and next-word prediction loss for LLaMA models in Figure 3F (r = 0.68, p = 0.21). We discuss our interpretations of these results and possible explanations in Section 6.2.
@@ 6.1 Implications for NLP: Building LLMs
Using brain alignment to interpret LLMs. Representations in the human brain support a wide range of downstream tasks that current models still struggle on. Prior works have used brain activity results to interpret neural networks (Dong & Toneva, 2023) and build more performant models (Dapello et al., 2020; 2022; Safarani et al., 2021). Instruction-tuning has emerged as a breakthrough technique to improve LLM abilities, quality of outputs, and allow LLMs to adapt to new tasks with minimal task-specific training. However, the manner in which instruction-tuning alters the internal representations of LLMs to achieve these improvements remains an open question. Brain activity data provides a neuroscientific angle to investigate this question. Our results show that instruction-tuning LLMs improves both world knowledge representations and human brain alignment. They suggest that instruction-tuning improves the ability of an LLM to retrieve world knowledge stored in its model weights and that the mechanisms that encode world knowledge in LLMs also improve representational alignment to the human brain. Overall, our results highlight the importance of integrating world knowledge in the development of future LLMs.
@@ 6.2 Implications for Neuroscience: Studying LLM-Human Alignment
Instruction-tuned LLMs are useful for studying LLM properties underlying brain and behavioral alignment.   To identify why LLM and human brains exhibit representational similarities, prior work has mostly focused on high-level properties such as model size (Antonello et al., 2023), and external behaviors such as predicting missing words Schrimpf et al. (2021); Caucheteux & King (2022). However, a key to understanding these similarities is to identify lower-level or internal properties of LLMs that underlie brain alignment. This includes the amount of knowledge LLMs contain, e.g., factual (AlKhamissi et al., 2022) and commonsense (Sap et al., 2020; Bosselut et al., 2019). Our work shows that we can harness instruction-tuned LLMs for this purpose as they have been trained to respond to a general question format, allowing us to evaluate LLMs in a more fine-grained manner. This allows the study of both internal (e.g., knowledge) and external (e.g., behavior) properties of LLMs, and how they correlate with brain and behavioral alignment.
Examining more dimensions of behavior.   To evaluate behavioral alignment, our work and many prior works compare LM and human next-word surprisal on reading tasks (Wilcox et al., 2020; Schrimpf et al., 2021; Eghbal A. Hosseini et al., 2023). This evaluates only a single dimension of LM and human behavior (per-word perplexity and reading times). On the models we test here, behavioral alignment is not significantly correlated with model size, world knowledge, or next-word prediction ability. While next-word prediction performance correlates with alignment to human reading times across a broad band of models (Schrimpf et al., 2021), this trend does not hold up in recent Transformer-based LMs (Oh & Schuler, 2023), having a surprising negative correlation with parameter count (Oh et al., 2022). Our results highlight the need to create more benchmarks to expand the dimensions of behavior examined for both LLMs and humans, in order to holistically evaluate LLM behavior, as well as LLM-human behavioral alignment.
Brain alignment datasets with humans performing diverse tasks.   Our work studies brain alignment using neural datasets that are limited to humans reading naturalistic stories and sentences in English. Unfortunately, there does not exist brain activity data for human participants attempting the BBH and MMLU benchmarks. This may explain why brain alignment is not significantly correlated with many categories of problem-solving on BBH, e.g., language understanding. In the future, we hope to study brain alignment with human participants performing more diverse sets of tasks, e.g., reading computer program code (Ivanova et al., 2020). This can identify more factors underlying LLM-brain alignment, and provide insights into how brain activity and the human language system may be shaped by various forms of problem-solving. Furthermore, for the neural datasets in our work, many of the larger models exceed the noise ceiling estimates (Appendix E), highlighting the need for more neural datasets (with better ways of computing noise ceiling estimates).
World knowledge shapes brain activity.   Our results show that world knowledge is a key factor in aligning LLM representations to human brain activity. LLMs with greater world knowledge across all tested subject domains produce representations that align more closely to human brain activity. Analogously, this suggests that world knowledge influences human brain activity, shaping the language comprehension systems in the brain.
@@ 6.3 Limitations and Future Work
Limited model families.   A model’s performance on MMLU and BBH is affected not only by the quantity of knowledge it possesses, but also by its ability to follow each benchmark’s instruction format. Hence, we restricted our study to two model families (T5 and LLaMA). This attempts to ensure that the models are trained on a similar instruction format, so they have similar ability to follow the instruction format of MMLU and BBH. Thus, any differences in their MMLU or BBH performance would point to differences in the quantity of knowledge they possess. However, we hope to add a greater variety of LLMs in future work to strengthen the generalizability of our claims.
Computational cost.   We investigate 25 LLMs, with the largest models having 33B parameters, on many datasets: brain alignment (Pereira2018, Blank2014, Wehbe2014), behavioral alignment (Futrell2018), next-word prediction (WikiText-2) and other evaluations (BBH, MMLU). Unfortunately, this makes it computationally expensive to replicate our experiments. In the hopes of addressing this concern, we provide our full results in Appendices E, F, G, H, and I.
@@ 7 Conclusion
We investigate whether instruction-tuning improves the alignment of LLMs to the human language system. We evaluate 25 LLMs with parameter sizes ranging from 77 million to 33 billion, across three neural datasets of humans reading naturalistic stories and sentences. We find that instruction-tuning generally improves the alignment of LLM representations to brain activity. Delving into the factors underlying LLM-brain alignment, we discover that world knowledge and model size are key determinants of brain alignment. This suggests that world knowledge helps shape representations in the human language system, and highlights the significance of integrating world knowledge in the development of future LLMs.
Reproducibility Statement
All models and code repositories utilized in this study are open-source and their corresponding links are provided in Appendix B and D for full transparency and reproducibility. For the calculation of brain and behavioral alignment, we relied on the Brain-Score repository (www.github.com/brain-score/language), which serves as a publicly accessible resource for conducting these assessments. We encourage researchers interested in replicating our findings to refer to the provided links and consult the Brain-Score repository for further details on datasets and the evaluation process. For measuring the performance of LLMs on BBH and MMLU benchmarks, we use the broadly-used instruct-eval repository (https://github.com/declare-lab/instruct-eval) with default settings.

### v2

@@ 4.2 Factors underlying LLM-brain alignment
To study why LLMs and human brains align in their representations, we compute the Pearson correlation between LLM brain alignment and various properties of LLMs: performance on a benchmark testing various reasoning abilities (BBH; Suzgun et al., 2022), performance on a benchmark testing world knowledge in various domains (MMLU; Hendrycks et al., 2021), language modeling ability, and model size.
@@
MMLU and BBH.   MMLU is designed to measure the world knowledge of LLMs across many subject domains. It contains 57 tasks, categorized into four world knowledge subject domains: STEM, Humanities, Social Sciences, and Others (a broad category ranging from finance to marketing to professional medicine). BBH contains 23 tasks, grouped into four types of problem-solving: Algorithmic and Multi-Step Arithmetic Reasoning; Natural Language Understanding; Use of World Knowledge; and Multilingual Knowledge and Reasoning. For both benchmarks, we follow the category classifications from the original papers. We perform the evaluations using the instruct-eval repository [^2] https://github.com/declare-lab/instruct-eval with default settings (3-shots for BBH, 5-shots for MMLU) and preset prompts. We measure the Pearson correlation (and its p-value) between LLM-brain alignment and performance in each category of MMLU and BBH. We obtain p-values after false discovery rate (FDR) correction.
@@
World Knowledge. We find that brain alignment is significantly and strongly correlated with world knowledge. Brain alignment is highly correlated with the MMLU Overall score (r = 0.81, p $<$ 0.001, Figure 2A), which reports the mean performance across all world knowledge subject domains on MMLU. Similarly, brain alignment is also strongly correlated with performance on tasks in the world knowledge category of BBH (r = 0.68, p $<$ 0.005; Figure 2B). Notably, brain alignment is poorly correlated with all other dimensions of BBH (see Table 2), though this could also be due to limitations of the tested models, as indicated by their low raw performance scores on some tasks. Overall, our results provide a strong signal that more accessible representations of world knowledge are a key factor in aligning LLM representations to human brain activity.
@@
Language Modeling Ability.   Prior works have shown correlations between brain alignment and next-word prediction (NWP) ability (Caucheteux & King, 2022; Schrimpf et al., 2021). We find similar results for correlation between brain alignment and NWP loss (r = -0.54, p $<$ 0.05, Appendix I). Interestingly, the strength of the correlation is weaker than that between brain alignment and world knowledge performance (r = 0.81). This suggests that world knowledge understanding is a better predictor of brain alignment than NWP ability.
@@
Model Size.   Finally, we find that brain alignment is significantly and strongly correlated with model size (r = 0.95, p $<$ 0.001, Figure 2C), as measured by the logarithm of the number of model parameters. Schrimpf et al. (2021) observe such a pattern for language models, and we find the pattern holds for instruction-tuned models, and models trained at a larger scale than their study (7B+ parameters). However, model size alone does not determine brain alignment. Our results show that smaller instruction-tuned LLMs can have greater brain alignment than larger vanilla models. For example, LLaMA-13B obtains brain alignment of 0.220, Vicuna-13B obtains 0.229, LLaMA-33B obtains 0.227, and Vicuna-33B obtains 0.232. Hence, Vicuna-13B has greater brain alignment than LLaMA-33B, due to instruction-tuning, despite being less than 40% its size. We observe a similar trend in another four models: T5-base, Flan-T5-base, T5-large, Flan-T5-large. Also, prior works showed that large random models achieve poor brain alignment (Schrimpf et al., 2021). These results demonstrate there are LLM properties aside from model size that contribute significantly to brain alignment.
@@
Neural datasets.   Brain alignment is strongly correlated with world knowledge and model size on Pereira2018 and Wehbe2014, but less strongly correlated on Blank2014, possibly because Blank2014 has fewer participants (N = 5) leading to greater noise in the results.
@@ 5 Behavioral Alignment
Figure 3: Instruction-tuning LLMs generally does not improve behavioral alignment to human reading times. Furthermore, behavioral alignment correlates poorly with all other tested measures: world knowledge, model size, and next-word prediction (NWP) ability. To compute behavioral alignment, we use the Futrell2018 benchmark in Brain-Score. The same language stimuli (naturalistic stories) are presented to LLMs and human participants. We then compute the Pearson correlation between per-word LLM perplexity and human reading times as the behavioral alignment. (A) Instruction-tuning does not generally improve behavioral alignment. Furthermore, behavioral alignment is poorly and not significantly correlated with all other measures: (B) world knowledge (p = 0.76), (C) model size (p = 0.31), (D) NWP loss for T5 models (p = 0.54), and (E) NWP loss for LLaMA models (p = 0.21). In Appendix M, we provide a larger, labeled version of this figure.
@@
In the previous section, we show that instruction-tuning aligns the internal representations of LLMs to human brain activity (Section 4.1). In this section, we explore whether instruction-tuning also aligns LLM behavior to human behavior.
@@
Following the approach of Schrimpf et al. (2021) implemented in the Brain-Score package (Schrimpf et al., 2020), we measure behavioral alignment by evaluating the similarity between LLM per-word perplexity and human per-word reading times given the same language stimuli (Appendix D). We use the self-paced reading times dataset from Futrell et al. (2018), consisting of the reading times of 179 human participants recorded while they were visually presented with 10 naturalistic stories. We provide language stimuli from this data as input to LLMs and evaluate behavioral alignment by computing the Pearson correlation between per-word LLM perplexity and per-word human reading times.
One potential concern in the behavioral alignment methodology is that human reading time is influenced by word length; longer words may take more time to read, even if they are not more challenging or surprising. Following Schrimpf et al. (2021), we address this by computing the surprisal of a multi-token word as the sum of the surprisals of its individual tokens. Importantly, we use consistent tokenizers when comparing vanilla LLMs against their instruction-tuned versions, thereby controlling for the effect of word length when evaluating the impact of instruction-tuning on behavioral alignment.
5.1 Instruction-tuning generally does not improve behavioral alignment
@@
Using the same models as in Section 4, we compare the behavioral alignment of each instruction-tuned LLM against its vanilla counterpart. Our results indicate instruction-tuning generally does not improve behavioral alignment to human reading times (Figure 3A). For half of the LLMs tested, it results in no change or reduced behavioral alignment. Our results align with recent studies (Gao et al., 2023; Steuer et al., 2023) that show improving LLM performance does not necessarily improve alignment with human reading behavior.
5.2 Factors underlying behavioral alignment
Next, we test the correlation between LLM behavioral alignment and model size, next-word prediction ability, various reasoning abilities (measured by BBH), and world knowledge across various domains (measured by MMLU). Contrary to our findings on the brain alignment correlations with model size and world knowledge (Section 4.2), we find that LLM behavioral alignment is not correlated with these factors: world knowledge (p = 0.76, Figure 3B), model size (p = 0.31, Figure 3C), next-word prediction loss for T5 models (p = 0.54, Figure 3D), and next-word prediction loss for LLaMA models (p = 0.21, Figure 3E).
@@ 6.1 Implications for NLP: Building LLMs
Using brain alignment for LLM interpretability. Prior works have used human brain activity to interpret models (Dong & Toneva, 2023) and build more performant models (Dapello et al., 2020; Safarani et al., 2021; Dapello et al., 2022). Instruction-tuning has emerged as a breakthrough technique to improve LLM abilities and quality of outputs, and allows LLMs to adapt to new tasks with minimal task-specific training. However, how instruction-tuning alters LLM internal representations to achieve these improvements remains an open question. Brain activity provides a neuroscientific angle to investigate this question. Our results show that instruction-tuning improves both performance on world knowledge benchmarks and human brain alignment, suggesting that the mechanisms that encode world knowledge in LLMs also align LLM representations to the human brain.
@@ 6.2 Implications for Neuroscience: Studying LLM-Human Alignment
Instruction-tuned LLMs are useful for studying LLM properties underlying brain and behavioral alignment.   To identify why LLM and human brains share representational similarities, prior work has mostly focused on high-level properties such as model size (Antonello et al., 2023), as well as external behaviors such as predicting missing words (Schrimpf et al., 2021; Caucheteux & King, 2022). However, a key to understanding these similarities is to identify internal properties of LLMs that underlie brain alignment, including the amount of knowledge LLMs encode, e.g., factual (AlKhamissi et al., 2022) and commonsense (Bosselut et al., 2019; Sap et al., 2020). Our work is the first to show that we can harness instruction-tuned LLMs for this purpose. Because they have been trained to respond to a general instruction format, we can evaluate LLMs on diverse tasks in a fine-grained manner, allowing the study of LLM properties both internal (e.g., knowledge) and external (e.g., behavior), and how they correlate with brain and behavioral alignment.
World knowledge shapes language comprehension and brain activity.   Our results show that world knowledge is a key factor in LLM brain alignment. LLMs demonstrating greater world knowledge across all tested subject domains have representations more similar to the human brain. This suggests that world knowledge influences human brain activity and shapes the language comprehension systems in our brain.
@@ 6.3 Limitations and Future Work
Examining more dimensions of behavior.   Our work and many prior works compare LLM and human next-word surprisal on reading tasks (Wilcox et al., 2020; Schrimpf et al., 2021; Eghbal A. Hosseini et al., 2023), which evaluates only a single behavioral dimension for LLMs (per-word perplexity) and humans (reading times). For the models we tested, behavioral alignment is not significantly correlated with model size, world knowledge, or next-word prediction ability. While next-word prediction performance correlates with alignment to human reading times across many LMs (Schrimpf et al., 2021), this trend does not hold up in recent transformer models (Oh & Schuler, 2023), having a surprising negative correlation with parameter count (Oh et al., 2022). In the future, we hope to evaluate LLMs on a wider range of behavioral dimensions (e.g., van Duijn et al., 2023; Koo et al., 2023; Kauf et al., 2024) to more holistically evaluate LLM-human behavioral alignment.
Brain alignment datasets with humans performing diverse tasks.   We study brain alignment using neural datasets of humans reading naturalistic stories and sentences in English. It would be interesting to study brain alignment to human participants attempting the BBH and MMLU benchmarks, but this data unfortunately does not exist. This may explain why brain alignment is not significantly correlated with many categories of problem-solving on BBH, e.g., algorithmic reasoning. In the future, we hope to study brain alignment with human participants performing more diverse sets of tasks, e.g., reading computer program code (Ivanova et al., 2020). This can identify more factors underlying LLM-brain alignment, and provide insights into how brain activity and the human language system may be shaped by various forms of problem-solving. Furthermore, some of the larger models exceed the noise ceiling estimates in our neural datasets (Appendix G), highlighting the need for more neural datasets and better ways of computing noise ceiling estimates.
@@ 7 Conclusion
We investigate whether instruction-tuning improves the alignment of LLMs to the human language system. We evaluate 25 LLMs with parameter sizes ranging from 77 million to 33 billion, across three neural datasets of humans reading naturalistic stories and sentences. We find that instruction-tuning generally improves the alignment of LLM representations to brain activity. Exploring the factors underlying LLM-brain alignment, we discover that world knowledge and model size are key determinants of brain alignment. This suggests that world knowledge helps shape representations in the human language system, and highlights the importance of integrating world knowledge in developing future LLMs.
8 Reproducibility Statement
We provide our full results in Appendices G, H, and L. We consider this important as our experiments may be computationally expensive to replicate. We investigated 25 LLMs, with the largest models having 33B parameters, on many datasets: brain alignment (Pereira2018, Blank2014, Wehbe2014), behavioral alignment (Futrell2018), next-word prediction (WikiText-2) and other evaluations (BBH, MMLU).
All models and code repositories used in our study are open-source and their corresponding links are provided in Appendix B and F for full transparency and reproducibility. To calculate brain and behavioral alignment, we used the Brain-Score repository (www.github.com/brain-score/language), a publicly accessible resource for conducting these assessments. We encourage researchers interested in replicating our findings to refer to the provided links and consult the Brain-Score repository for further details on datasets and the evaluation process. To evaluate LLMs on BBH and MMLU, we utilize the widely-used instruct-eval repository (https://github.com/declare-lab/instruct-eval) with default settings.